Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

probability

Expert replies
by rahul.s » Sun Jan 31, 2010 3:33 am
A jar contains equal amounts of green jellybeans and red jellybeans. Mark selects at random one jellybean at a time from the jar, notes its color, and returns it to the jar. If he repeats this process four times, what is the probability that he selects 3 green jellybeans and 1 red jellybean?

A) 1/4
B) 3/16
C) 1/8
D) 1/16
E) 1/64

OA: A
Last edited by rahul.s on Sun Jan 31, 2010 5:01 am, edited 1 time in total.
Join the discussion
Source: — Problem Solving |

by ajith » Sun Jan 31, 2010 3:44 am
rahul.s wrote:A jar contains equal amounts of green jellybeans and red jellybeans. Mark selects at random one jellybean at a time from the jar, notes its color, and returns it to the jar. If he repeats this process four times, what is the probability that he selects 3 green jellybeans and 1 red jellybean?

A) 1/4
B) 3/16
C) 1/8
D) 1/16
E) 1/64

OA: A

[spoiler]my solution: i assumed 4 jellybeans in each jar. we need 3 green and 1 red = (1/2 + 1/2 +1/2) * 1/2 = 3/6 * 1/2 = 1/4. but the OE was different. is it correct?[/spoiler]
Total no of outcomes = 2^4

Favorable outcomes = 4 ( RRRG, GRRR,RRGR, RGRR)

Probability = 4/16 =1/4
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by rahul.s » Sun Jan 31, 2010 3:54 am
ajith,

could you please simplify that for me? why is it 2^4? why are the the favorable outcomes = 4? where did i err?
Join the discussion

by ajith » Sun Jan 31, 2010 3:58 am
rahul.s wrote:ajith,

could you please simplify that for me? why is it 2^4? why are the the favorable outcomes = 4? where did i err?
From The bag you can either take out a Red one or a green one Denoted By R or G

Now u can make the first choice in 2 ways, second in 2 , third in 2 and the fourth in 2 (either R or G)

hence total events = 16

Now favourable events are ( RRRG, GRRR,RRGR, RGRR) out of all these
Always borrow money from a pessimist, he doesn't expect to be paid back.
Join the discussion

by bhumika.k.shah » Sun Jan 31, 2010 5:01 am
Rahul how is 1/2 + 1/2 + 1/2 = 3/6??
shouldnt it be 3/2??
rahul.s wrote:A jar contains equal amounts of green jellybeans and red jellybeans. Mark selects at random one jellybean at a time from the jar, notes its color, and returns it to the jar. If he repeats this process four times, what is the probability that he selects 3 green jellybeans and 1 red jellybean?

A) 1/4
B) 3/16
C) 1/8
D) 1/16
E) 1/64

OA: A

[spoiler]my solution: i assumed 4 jellybeans in each jar. we need 3 green and 1 red = (1/2 + 1/2 +1/2) * 1/2 = 3/6 * 1/2 = 1/4. but the OE was different. is it correct?[/spoiler]
Join the discussion

by bhumika.k.shah » Sun Jan 31, 2010 5:16 am
Hey ,
can anyone explain the formula way ?? if there are 4 green colored jellybeans and 4 red colored jelly beans????
Join the discussion

by viidyasagar » Wed Sep 15, 2010 6:17 pm
bhumika.k.shah wrote:Hey ,
can anyone explain the formula way ?? if there are 4 green colored jellybeans and 4 red colored jelly beans????
I used Stuart Kovinsky's coin flip approach to solve this question in less than 30 seconds!!!

the real challenge, though, is to figure that this question indeed merits a coin flip solution


https://www.beatthegmat.com/coin-flip-qu ... html#75414

The link also contains a formula based approach!!!
Join the discussion